# Seeing the Unseen: Visual Metaphor Captioning for Videos

### v1

@@ Abstract
Metaphors are a common communication tool used in our day-to-day life. The detection and generation of metaphors in textual form have been studied extensively but metaphors in other forms have been under-explored. Recent studies have shown that Vision-Language (VL) models cannot understand visual metaphors in memes and adverts. As of now, no probing studies have been done that involve complex language phenomena like metaphors with videos. Hence, we introduce a new VL task of describing the metaphors present in the videos in our work. To facilitate this novel task, we construct and release a manually created dataset with $\mathbf{705}$ videos and $\mathbf{2115}$ human-written captions, along with a new metric called Average Concept Distance (ACD), to automatically evaluate the creativity of the metaphors generated. We also propose a novel low-resource video metaphor captioning system: GIT-LLaVA, which obtains comparable performance to SoTA video language models on the proposed task. We perform a comprehensive analysis of existing video language models on this task and publish our dataset, models, and benchmark results to enable further research.
Seeing the Unseen: Visual Metaphor Captioning for Videos
Abisek Rajakumar Kalarani , Pushpak Bhattacharyya , Sumit Shekhar Department of Computer Science and Engineering, IIT Bombay, India Adobe Research, India {abisekrk, pb}@cse.iitb.ac.in sushekha@adobe.com
@@ 1 Introduction
Figurative languages in textual form have been well-studied in literature Abulaish et al. (2020). With the advent of powerful AI assistants like ChatGPT and BARD and tools that are built on top of them, it is possible to interact with these AI systems through images and audio. Hence it becomes important to build and test models that can handle complex language phenomena, such as metaphors, across multiple modalities. Recent works on Visual metaphors (Yosef et al. 2023, Chakrabarty et al. 2023) focus on understanding metaphors present in images and generating images from prompts with metaphors. They show that it is challenging to deal with metaphors presented visually.
@@
Recently, chat assistants that can answer questions related to videos have shown good promise on standard video datasets (Zhang et al. 2023; Li et al. 2023b; Maaz et al. 2023). However, they struggle to understand videos that contain metaphors. To this effect, we introduce a novel task of ‘Video Metaphor Captioning’ (VMC) that involves describing the metaphors present in the video with a single-line caption. We manually annotate a dataset of videos with metaphor information. We also build and release a novel low-resource video metaphor captioning model that achives comparable performance to SOTA video language models on video metaphor captioning task despite being trained on limited pretraining data.
@@
    A novel low-resource Vision-Language model (GIT video model followed by Vicuna LLM) pretrained and fine-tuned for video metaphor understanding (Section 4).
@@ 2 Related Work
Metaphor Detection is the task of classifying if the given sentence/token contains a metaphor or not. In recent years, metaphor detection has been explored with the aid of large language models. Choi et al. (2021) used the contextual embeddings from BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) with a late interaction mechanism to make use of linguistic metaphor identification theories. Aghazadeh et al. (2022) probed and analyzed the metaphorical language encoded in the large language models. Su et al. (2020) used a combination of global sentence features and POS information to perform token-level metaphor detection. Badathala et al. (2023) used a multitasking approach to detect hyperbole and metaphors together.
@@
Visual Metaphors: The detection and generation of metaphors in textual form have been explored extensively but the use of metaphors in other modalities like images is not explored until very recently. Akula et al. (2022) introduced a set of tasks related to understanding visual metaphors. They showed that existing Vision-Language models are not good at understanding visual metaphors. Yosef et al. (2023) introduced a multimodal dataset that contains metaphors, similes, and idioms with corresponding images for them. Zhang et al. 2021, Hwang and Shwartz 2023, and Xu et al. 2022 explored the uses of metaphors in memes and released datasets for understanding metaphors in memes. Chakrabarty et al. (2023) explored generating visual metaphors from metaphorical input sentences. They release a dataset called HAIVMet which contains 6476 images of visual metaphors generated with DALL-E 2 Ramesh et al. (2022).
@@ 3.1 Annotation Details
Figure 3: An overview of our Video Metaphor Captioning system, GIT-LLaVA. The text decoder representation of GIT is mapped to the embedding space of Vicuna to generate metaphor captions.
@@ 3.2 Dataset Statistics
All videos are accompanied by three captions. Our Video Metaphor Captioning Dataset (VMCD) consists of $705$ metaphoric videos with $2115$ captions. The train, validation, and test split contain $400$, $55$, and $250$ videos each with $1200$, $165$, and $750$ captions respectively.
@@ 4 Our Model
We use the LLaVA-13B-V1.5 Liu et al. (2023) model architecture for our experiments inspired by its success on many Vision Language tasks. We use the Generative Image Text Transformer model (GIT) Wang et al. (2022) as the video captioning model (C) for obtaining the video representation and Vicuna Zheng et al. (2023) as the LLM. In all our experiments we freeze the weights of the GIT model and only finetune the mapping network and the LLM. Since we train the mapping network to learn the mapping between the GIT decoder state to the embedding space of the LLM, the mapping network maps GIT’s understanding of the video in the form of its representation to the LLM’s embedding space, allowing the LLM to directly generate output from the video. This also reduces the need to pretrain the model on a huge corpus of Video-Text parallel data which is resource intesive.
@@ 5.1 Pretraining
Our video metaphor captioning system uses a pretrained video captioning model to obtain video representation. The video representation needs to be mapped to the embedding space of the LLM for it to generate fluent captions. Our VMC dataset is small and may not be sufficient to learn this mapping. Hence, we initially pretrain the model on pretraining dataset of images.
@@ 5.2 Video Metaphor Captioning
Table 1: Experimental results on our VMC dataset in comparison to other models. ACS and ACD denote the Average Concept Similarity and Average Concept Distance metric weighted by BERTScore respectively. Cosine similarity and distance is computed between the concepts compared in the metaphor caption. The best model is in bold and the next-best model is underlined.
6 Evaluation Metrics
[…]
S-BERT Reimers and Gurevych (2019) (all-mpnet-base-v2) is used to obtain representations for PC and SC. For captions that do not contain either PC or SC, the similarity score is set as 1 to penalize the model. Thus the model is evaluated based on how diverse comparison it can make for the object in question. We also discuss the correlation between the proposed ACD metric and human evaluation of metaphor creativity in Section A.6. In addition to these automated metrics, we also manually evaluate models on four metrics- Fluency, Primary Concept Consistency, Consistency, and Creativity.
@@ 7.2 Human Evaluation
In addition to automated metrics, we also perform human evaluation (Figure 4) on 100 videos from the test set with outputs generated by all models. Three annotators in the age group of 25-30 were used to annotate these videos. Each annotator annotated 50 videos. 25 videos were common among the annotators. The annotation was done on four metrics- Fluency, Primary Concept Consistency, Consistency, and Creativity. These metrics are discussed in detail along with Inter Annotator Agreement (IAA) in Appendix A.7.
@@ A.3 Prompts for Training
A.4 Dataset Statistics
VMC dataset consists of $705$ videos with $2115$ captions. The average duration of the video is $54$ seconds, and the average length of the caption is $8.9$ words. Figure 6 shows histograms for the distribution of video duration and caption lengths.
A.5 Ablation Study
@@ A.5.2 Synthetic Dataset
We use images and captions from the MSCOCO dataset Lin et al. (2014). We prompt GPT-3.5-turbo model with the following prompt: “Convert the following image caption to a metaphoric image caption in the following format <primary concept> is as <property> as <secondary concept>. Input: mscoco_caption”. For example, we convert the image caption ‘A bicycle replica with a clock as the front wheel’ to ‘A timepiece is as cyclical as a bicycle’s revolution’. The generated captions were then cleaned to remove captions that did not follow the template in the prompt. The final pretraining dataset consists of $90886$ images and corresponding synthetically generated metaphoric captions. We further evaluate the quality of the generation by manual evaluation. We employed two annotators to annotate the fluency and creativity of the generated captions. The annotators provided binary classification labels for fluency and creativity. The captions were $98.7$% fluent and $97.8$% creative. This confirms the quality of the synthetic data.
Rows 3 and 4 in Table 2 show results of the GIT-LLaVA and the GIT-LLaVA-X models trained on this synthetic data. The results are comparable to the models trained on the previously discussed pretraining data. No improvement is performance was observed.
A.5.3 No Pretraining
@@
A.5.4 Additional Video Components
@@
A.6 Average Concept Distance Metric
We compute the correlation of the Average Concept Distance (ACD) metric with the human evaluation of captions. The manual evaluation in Section 7.2 was done for 100 videos with 5 captions for each video. We used the ACD metric evaluator to compare the ACD scores with binary labels provided by the annotators. The ACD scores and binary labels had a Pearson correlation coefficient of $0.403$ with p-value $<<0.0001$. As creativity is a subjective metric, the moderate correlation is very significant.
A.7 Manul Evaluation Results
The manual evaluation was done on four metrics- Fluency, Creativity, Primary Concept Consistency, and Consistency by three annotators on 100 videos. Each annotator gave binary labels for each generated caption on all 4 metrics.
The manual evaluation scores further complement the results obtained with automatic metrics. Table 3 shows the Inter-Annotator Agreement (IAA) between the three annotators for manual evaluation of captions generated by all 6 models. The scores indicate substantial agreement between the reviewers. Both set of annotators used for annotating VMCD and the manual evaluations received fair and competitive stipends.
Cohen’s Kappa ($\kappa$) A B
B 0.639
C 0.637 0.685
Fleiss’ Kappa (K) 0.653
@@ A.7 Manul Evaluation Results
A.8 Examples
Figure 7 shows some examples of wrong captions generated by the models.

### v2

# Unveiling the Invisible: Captioning Videos with Metaphors

@@ Abstract
Metaphors are a common communication tool used in our day-to-day life. The detection and generation of metaphors in textual form have been studied extensively but metaphors in other forms have been under-explored. Recent studies have shown that Vision-Language (VL) models cannot understand visual metaphors in memes and adverts. As of now, no probing studies have been done that involve complex language phenomena like metaphors with videos. Hence, we introduce a new VL task of describing the metaphors present in the videos in our work. To facilitate this novel task, we construct and release a manually created dataset with $\mathbf{705}$ videos and $\mathbf{2115}$ human-written captions, along with a new metric called Average Concept Distance (ACD), to automatically evaluate the creativity of the metaphors generated. We also propose a novel low-resource video metaphor captioning system: GIT-LLaVA, which obtains comparable performance to SoTA video language models on the proposed task. We perform a comprehensive analysis of existing video language models on this task and publish our dataset, models, and benchmark results to enable further research. [^1] Code, Data, and Models are available at https://github.com/abisekrk/video-metaphor-captioning
Unveiling the Invisible: Captioning Videos with Metaphors
Abisek Rajakumar Kalarani , Pushpak Bhattacharyya , Sumit Shekhar Department of Computer Science and Engineering, IIT Bombay, India Adobe Systems, India {abisekrk, pb}@cse.iitb.ac.in sushekha@adobe.com
@@ 1 Introduction
Motivation: Figurative languages in textual form have been well-studied in literature Abulaish et al. (2020). With the advent of powerful AI assistants like ChatGPT and BARD and tools that are built on top of them, it is possible to interact with these AI systems through images and audio. Hence it becomes important to build models that can handle complex language phenomena, such as metaphors, across multiple modalities. Recent works on Visual metaphors (Yosef et al. 2023, Chakrabarty et al. 2023) focus on understanding metaphors present in images and generating images from prompts with metaphors. They show that it is challenging to deal with metaphors presented visually.
@@
Recently, chat assistants that can answer questions related to videos have shown good promise on standard video datasets (Zhang et al. 2023; Li et al. 2023b; Maaz et al. 2023). However, they struggle to understand videos that contain metaphors. To this effect, we introduce a novel task of ‘Video Metaphor Captioning’ (VMC) that involves describing the metaphors present in the video with a single-line caption. We manually annotate a dataset of videos with metaphor information. We build and release a novel low-resource video metaphor captioning model that achives comparable performance to SOTA video language models on video metaphor captioning task despite being trained on limited pretraining data.
@@
    A novel low-resource Vision-Language model (GIT video model followed by Vicuna LLM) pretrained and fine-tuned for video metaphor captioning (Section 4).
@@ 2 Related Work
Metaphor Detection is the task of classifying if the given sentence/token contains a metaphor or not. In recent years, metaphor detection has been explored with the aid of large language models. Choi et al. (2021) used the contextual embeddings from BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) with a late interaction mechanism to make use of linguistic metaphor identification theories. Aghazadeh et al. (2022) probed and analyzed the metaphorical language encoded in the large language models. Su et al. (2020) used both global sentence features and POS information to perform token-level metaphor detection. Badathala et al. (2023) used a multitasking approach to detect hyperbole and metaphors together.
@@
Visual Metaphors: The detection and generation of metaphors in textual form have been explored extensively but the use of metaphors in other modalities like images is not explored until very recently. Akula et al. (2022) introduced a set of tasks related to understanding visual metaphors. They showed that existing Vision-Language models are not good at understanding visual metaphors. Yosef et al. (2023) introduced a multimodal dataset that contains metaphors, similes, and idioms with corresponding images for them. Zhang et al. 2021, Hwang and Shwartz 2023, and Xu et al. 2022 explored the uses of metaphors in memes and released datasets for understanding metaphors in memes. Chakrabarty et al. (2023) explored generating visual metaphors from metaphorical input sentences. They release a dataset called HAIVMet which contains 6476 images of visual metaphors generated with DALL-E 2 Ramesh et al. (2022). Achlioptas et al. (2021) and Achlioptas et al. (2022) explore emotions invoked by images.
@@ 3.1 Annotation Details
Figure 3: An overview of our Video Metaphor Captioning system, GIT-LLaVA. In the pretraining stage, only the mapping network is trained with image datasets. In the finetuning stage, the text decoder representation of the GIT model is mapped to the embedding space of the Vicuna model to generate metaphor captions.
@@ 3.2 Dataset Statistics
All videos are accompanied by three captions. Our Video Metaphor Captioning Dataset (VMCD) consists of $705$ metaphoric videos with $2115$ captions. The train, validation, and test split contain $400$, $55$, and $250$ videos each with $1200$, $165$, and $750$ captions respectively. The average duration of the video is $54$ seconds, and the average length of the caption is $8.9$ words. Figure 7 shows histograms for the distribution of video duration and caption lengths.
@@ 4 Our Model
We use the LLaVA-13B-V1.5 Liu et al. (2023) model architecture for our experiments inspired by its success on many Vision Language tasks. We use the Generative Image Text Transformer model (GIT) Wang et al. (2022) as the video captioning model (C) for obtaining the video representation and Vicuna Zheng et al. (2023) as the LLM. In all our experiments we freeze the weights of the GIT model and only finetune the mapping network and the LLM. Our model has two key advantages. Since we train the mapping network to learn the mapping between the GIT decoder state to the embedding space of the LLM, the mapping network maps GIT’s understanding of the video in the form of its representation to the LLM’s embedding space, allowing the LLM to directly generate output from the video. This also reduces the need to pretrain the model on a huge corpus of Video-Text parallel data which is resource intensive.
@@ 5.1 Pretraining
Our video metaphor captioning system uses a pretrained video captioning model to obtain video representation. The video representation needs to be mapped to the embedding space of the LLM for it to generate fluent captions. Our VMC dataset is small and may not be sufficient to learn this mapping. Hence, we initially pretrain the model on a pretraining dataset of metaphor and normal images, since no existing dataset has videos with metaphor information.
@@ 5.2 Video Metaphor Captioning
Figure 4: An overview of the “Average Concept Distance” metric. We compute the cosine distance between the primary and the secondary concept and ground it with the BERTScore.
6 Evaluation Metrics
[…]
S-BERT Reimers and Gurevych (2019) (all-mpnet-base-v2) is used to obtain representations for PC and SC. For captions that do not contain either PC or SC, the similarity score is set as 1 to penalize the model. Thus the model is evaluated based on how diverse comparison it can make for the object in question. Figure 4 provides an overview of the ACD metric computation. In addition to these automated metrics, we also manually evaluate models on four metrics- Fluency, Primary Concept Consistency, Consistency, and Creativity.
@@ 6 Evaluation Metrics
Table 1: Experimental results on our VMC dataset in comparison to other models. ACS and ACD denote the Average Concept Similarity and Average Concept Distance metric weighted by BERTScore respectively. Cosine similarity and distance is computed between the concepts compared in the metaphor caption. The best model is in bold and the next-best model is underlined. All reported scores are the mean scores of three runs.
@@ 7 Results and Analysis
We evaluate the models based on both automatic metrics and human evaluation.
@@ 7.2 Human Evaluation
In addition to automated metrics, we also perform human evaluation (Figure 5) on 100 videos from the test set with outputs generated by all models. Three annotators in the age group of 25-30 were used to annotate these videos. Each annotator annotated 50 videos. 25 videos were common among the annotators. The manual evaluation was done on four metrics by three annotators on 100 videos. Each annotator gave binary labels for each generated caption on all 4 metrics.
Pairwise Comparison Cohen’s Kappa ($\kappa$)
A vs B 0.639
B vs C 0.685
A vs C 0.637
Overall Fleiss’ Kappa (K) 0.653
The manual evaluation scores further complement the results obtained with automatic metrics. Table 2 shows the Inter-Annotator Agreement (IAA) between the three annotators for manual evaluation of captions generated by all 6 models. The scores indicate substantial agreement between the reviewers. Both sets of annotators used for annotating VMCD and the manual evaluations received fair and competitive stipends.
@@
We also compute the correlation of the Average Concept Distance (ACD) metric with the human evaluation of captions. We compared the ACD metric scores computed by our module with the binary labels provided by our annotators for the above 100 videos. The ACD scores and binary labels had a Pearson correlation coefficient of $0.403$ with p-value $<<0.0001$. As creativity is a subjective metric, the moderate correlation is very significant.
@@ A.3 Prompts for Training
A.4 Ablation Study
@@ A.4.2 Synthetic Dataset
We use images and captions from the MSCOCO dataset Lin et al. (2014). We prompt GPT-3.5-turbo model with the following prompt: “Convert the following image caption to a metaphoric image caption in the following format <primary concept> is as <property> as <secondary concept>. Input: mscoco_caption”. For example, we convert the image caption ‘A bicycle replica with a clock as the front wheel’ to ‘A timepiece is as cyclical as a bicycle’s revolution’. It is important to note that the MSCOCO dataset does not contain metaphoric images in them. We generate metaphoric captions for normal images, as the goal is only to learn the mapping between literal captions and metaphoric captions.
The generated captions were then cleaned to remove captions that did not follow the template in the prompt. The final pretraining dataset consists of $90886$ images and corresponding synthetically generated metaphoric captions. We further evaluate the quality of the generation by manual evaluation. We employed two annotators to annotate the fluency, creativity and consistency of a randomly sampled 500 generated captions. The annotators provided binary classification labels. The captions were $98.7$% fluent, $97.8$% creative, and $96.4\%$ consistent. This confirms the quality of the synthetic data generated by the model.
Rows 3 and 4 in Table 3 show results of the GIT-LLaVA and the GIT-LLaVA-X models trained on this synthetic data. The results are comparable to the models trained on the previously discussed pretraining data. No improvement in the performance was observed. Since the pretraining data had metaphoric captions to normal images, it leads to increase in hallucinations during finetuning stage as the model generated metaphors about things not present in the videos.
A.4.3 No Pretraining
@@
A.4.4 Additional Video Components